OpenStudy (anonymous):

an arithmetic sequence has a1 = a5 and a5 = 13. how many terms of this sequence must be added to obtain 60?

myininaya (myininaya):

if a_1=a_5 and a_5=13 then a_1=13 so sounds like your common difference is 0

myininaya (myininaya):

I'm afraid you will never be able to get exactly 60 by adding a some of the terms of the sequence

myininaya (myininaya):

mainly because 13 doesn't divide 60

myininaya (myininaya):

unless there is a mistake in the question

myininaya (myininaya):

maybe it wants us to get to the number n such that the sum is at least 60 a1+a2+a3+...+an=60 and that case we could do it so you have 13+13+13+...+13=13n solve 13n=60 for n

OpenStudy (anonymous):

I meant a_1=5 and a_5=13

myininaya (myininaya):

you can find the common difference by doing \[\frac{a_5-a_1}{5-1}\]

myininaya (myininaya):

let me know when you think you have the common difference

OpenStudy (anonymous):

2?

myininaya (myininaya):

right

myininaya (myininaya):

\[\text{ the sequence looks like } \\ 5,7,9,11,13,15,17,19,...\]

myininaya (myininaya):

5+7+9+11+13+15=?

myininaya (myininaya):

there is another way to do this if you don't like that

OpenStudy (anonymous):

so a_6?

myininaya (myininaya):

well it asked for how many terms we need to add

myininaya (myininaya):

so close you needed 6 of the first terms

OpenStudy (anonymous):

so 6 terms?

myininaya (myininaya):

just adding the terms til you get to 60

myininaya (myininaya):

\[Sum=\frac{n (a_1+a_n)}{2} \\ Sum=\frac{n(a_1+(a_1+(n-1)d))}{2} \\ \text{ you said } d=2 \text{ and } a_1=5 \text{ and we want sum=60 }\] \[60=\frac{n(5+(5+(n-1)2))}{2} \\ \text{ \times 2 on both sides } \\ 120=n(5+5+(n-1)2) \\ 120=n(10+2n-2) \\ 120=n(8+2n) \\ 120=2n^2+8n \\ \\ 2n^2+8n-120=0 \\ n^2+4n-60=0 \\ (n-6)(n+10)=0\]

myininaya (myininaya):

but you also get n=6 from this way n=-10 by the way doesn't make sense

myininaya (myininaya):

Anyways the way that uses a quadratic equation might be better for bigger numbers

myininaya (myininaya):

I will make it look a little pretty the formula that is that I used \[Sum=\frac{n(2a_1+(n-1)d)}{2} \\ Sum=\frac{n(2a_1+dn-d)}{2} \\ Sum=\frac{2 a_1 n +dn^2-dn}{2} \\ 2 Sum=2a_1 n+dn^2-dn \\ 0=2a_1n+dn^2-dn-2 Sum \\ 0=dn^2+(-d+2a_1 )n -2 Sum\]

myininaya (myininaya):

so you can use that bottom quadratic to find n

myininaya (myininaya):

should be factorable if there is an integer such that allows you to add a certain amount of terms together to get a certain sum

myininaya (myininaya):

\[n=\frac{d-2a_1 \pm \sqrt{(-d+2a_1)^2-4(d)(-2Sum)}}{2d} \\ n=\frac{d-2a_1 \pm \sqrt{(-d+2a_1)^2+8d Sum}}{2d}\] but if you don't like factoring you could use the quadratic formula as I did here one of the answers will not work just like I have shown about with the n=-10 which did not make sense :)

OpenStudy (anonymous):

ok thank you i think i got it now

myininaya (myininaya):

but hey I know that probably looks a little confusing pretend we have the arithemetic sequence 7,10,13,16,... common difference d=3 a_1=7 pretend we want sum 30 plug in where have that formula for n

OpenStudy (anonymous):

whoa nelly is this the problem \(a_1=a_5\)?

OpenStudy (anonymous):

of it is arithmetic, and \(a_1=a5\) then all the terms must be the same right?

myininaya (myininaya):

\[n=\frac{3-2(7) \pm \sqrt{(-3+2(7))^2+8(3)(30)}}{2(3)} \\ n=\frac{-11 \pm \sqrt{(11)^2+720}}{6} \\ \text{ going \to throw out the obvious negative value } \\ n=\frac{-11 + \sqrt{121+720}}{6} \\ n=\frac{-11 + \sqrt{841}}{6}\] \[n=\frac{-11+29}{6}=\frac{18}{6}=3\]

myininaya (myininaya):

which is true since 17+10+13=30

myininaya (myininaya):

anyways yeah he said that was a typeo earlier

OpenStudy (anonymous):

oh i came late

OpenStudy (anonymous):

\[5+7+9+11+13+...=60\]

myininaya (myininaya):

i prefer that way of adding up the terms til you get to 60

OpenStudy (anonymous):

\[\sum_{k=1}^n2k-1=n^2\]

myininaya (myininaya):

but if the sum you want is large you probably have to do it the icky way

OpenStudy (anonymous):

you have \(60\) and are missing \(1\) and \(3\) so you have 8 terms

myininaya (myininaya):

oh you want to do it that way thats cute too

OpenStudy (anonymous):

you are cute as a unicorn

myininaya (myininaya):

i wouldn't know anything about unicorns

OpenStudy (anonymous):

https://www.youtube.com/watch?v=oDFiri7HH1E